is this a solution of the equation 2(x-2)-5x=5(2-X);7
Answer by rapaljer(4671)

I think instead of solving for x, they want you to just substitute the value of 7 in place of all the x's. If it comes out a true statement, then it's a solution. If it is a false statement, then it is NOT a solution.

2(x -2) - 5x = 5(2-x)
2(7-2) -5(7) = 5(2-7)
2(5) - 35 = 5(-5)
10 - 35 = -25

Since this is a true statement, it is a solution!

R^2 at SCC